40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 22 18:55:31 2022, yuta, Update, LSC, Daily alignment work; POY trouble solved Screenshot_2022-06-22_17-17-42_XYaligned.pngScreenshot_2022-06-22_18-58-26_Transmission.png
    Reply  Wed Jun 22 19:41:13 2022, Koji, Update, LSC, Daily alignment work; POY trouble solved 
Message ID: 16940     Entry time: Wed Jun 22 18:55:31 2022     Reply to this: 16941
Author: yuta 
Type: Update 
Category: LSC 
Subject: Daily alignment work; POY trouble solved 

[Koji, Yuta]

I found that Yarm cannot be locked today. Both POY11 and POYDC were not there when Yarm was aligned, and ITMY needed to be highly misaligned to get POYDC.
POY beam also could not be found at ITMY table.
Koji suggested to use AS55 instead to lock Yarm. We did it (AS55_I_ERR, C1:LSC-YARM_GAIN=-0.002) and manually ASS-ed to get Yarm aligned (ASS with AS55 somehow didn't work).
After that, we checked ITMY table and found that POY beam was clipped at an iris which was closed!
I opened it and now Yarm locks with POY11 again. ASS works.
PMC was also aligned.

C1:PSL-PMC_PMCTRANSPD ~0.74
C1:IOO-MC_TRANS_SUM ~14000
C1:LSC-TRY_OUT ~0.7
C1:LSC-TRX_OUT ~0.8

Attachment 1: Screenshot_2022-06-22_17-17-42_XYaligned.png  56 kB  Uploaded Wed Jun 22 20:03:33 2022  | Show | Hide all | Show all
Attachment 2: Screenshot_2022-06-22_18-58-26_Transmission.png  34 kB  Uploaded Wed Jun 22 20:03:41 2022  | Hide | Hide all | Show all
Screenshot_2022-06-22_18-58-26_Transmission.png
ELOG V3.1.3-